OpenStudy (anonymous):

Angie was working on solving the exponential equation 23x = 6; however, she is not quite sure where to start. Using complete sentences, describe to Angie how to solve this equation and how solving would be different if the bases were equal.

OpenStudy (abmon98):

\[23^x=6\] Log both sides we dont use log base we use log base 23 because the inverse property of log base 23 of 23 raised to the power of x just equals x \[\log _{23}23^x=\log _{23}6\] \[x=\log_{6} /\log_{23} \]

OpenStudy (anonymous):

is that log sub 22 or 23 then 23^x

OpenStudy (abmon98):

log base 6 /log base 23

OpenStudy (anonymous):

in the second part of solving you have logsub something(22 or 23) then base 23^x=log(either 22 or23) 6

OpenStudy (abmon98):

i made a mistake, \[\log _{a}(x ^{b)}\]=\[blog _{a}x\] xlog(23)=log(6) Divide both sides by log(23)